OpenStudy (anonymous):

How do I get from 2a+a+sqrt(3a^2)=20 to.... a=8.453?

OpenStudy (anonymous):

How did you arrive at that answer?

OpenStudy (anonymous):

wolframalpha.com

OpenStudy (anonymous):

are you sure you plugged it in correctly?

OpenStudy (anonymous):

I entered the equation exactly how it here.

OpenStudy (anonymous):

Then you must've typed it here wrong. That is not the correct answer.

OpenStudy (anonymous):

Double check

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2a%2Ba%2B+sqrt%283a%5E2%29%3D20 are you able to see this, my friend?

OpenStudy (anonymous):

\[2a+a+\sqrt(3a^2)=20\]

OpenStudy (anonymous):

Yes, that correct.

OpenStudy (anonymous):

It doesn't say the answer is 8.453

OpenStudy (anonymous):

It does, in the solution: a=-10/3(sqrt(3)-3).

OpenStudy (anonymous):

That comes out to approximately 4.2265

OpenStudy (anonymous):

Not 8.453

OpenStudy (anonymous):

You're right I made a mistake, my answer is two times 4.2265.

OpenStudy (anonymous):

OK i will show you how to do the work correctly.

OpenStudy (anonymous):

I am still wondering how I get there though.

OpenStudy (anonymous):

\[3a+\sqrt(3a^2)=20\] \[a(3+\sqrt(3))=20\] \[a=\frac{20}{(3+\sqrt(3))}\]

OpenStudy (anonymous):

From here we rationalize the denominator

OpenStudy (anonymous):

I am confused in the second step.

OpenStudy (anonymous):

\[\frac{20}{3+\sqrt(3)} * \frac{3-\sqrt(3)}{3-\sqrt(3)}\]

OpenStudy (anonymous):

In the second step i simply factored out an a because \[\sqrt(a^2)=a\]

OpenStudy (anonymous):

so it would be \[3a+a \sqrt(3)\]

OpenStudy (anonymous):

Then you can factor out an a here, see?

OpenStudy (anonymous):

okay, i see now thank you very much !

OpenStudy (anonymous):

Do you know how to rationalize the denominator?

OpenStudy (anonymous):

You must do that to get to the final answer

OpenStudy (anonymous):

I do, i need to try it to make sure i get the mechanics of it.

OpenStudy (anonymous):

The numberator should be 60-20sqrt(3), correct? And I am having difficulties with the denominator? Could you help please?

OpenStudy (anonymous):

Nevermind, I figured it out, thank you.